

********************************************************************
UARS WEEKLY STATUS REPORT - 24 April 1998
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
***********
SUSIM
***********
Nothing received.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
HALOE continues successful operations with all engineering parameters
within nominal operating range and the science data results are of
excellent quality.
***********
MLS
***********
MLS was reactivated successfully on Apr 15, in the reduced 85 watt power
mode. MLS instrument operations have continued flawlessly. MLS was powered
off at 109::22:47 UT. MLS is scheduled to resume north viewing operations
starting Apr 29.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF will be unavailable early Friday afternoon April 24 - Monday morning
April 27 due to Building 23 power outage.
The CDHF has received a total of 386 CD-ROM requests as of 24-April-1998.
There were five orders received and five orders shipped.
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
25 days PEM (3.32) PROD (1379R-1149C=230L)
9 days CLAES (7.0) ENG (72R-60C-1F=11L)
711 days ATT (7.2) PROD (2178R-1914C=264L)
CDHF Pending software upgrades: IDL 5.1, OpenIngres, UCSS.
***********
FOT
***********
REPORTING PERIOD
This report covers 10 April 1998 (Orbit 35960, GMT Day 101) through 17
April 1998 (Orbit 36063, GMT Day 107).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and
is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
SUSIM 104/2334 36019 Off
ACRIM 104/2334 36019 Off
HRDI 104/2334 36019 Off
MLS 105/0015 36019 On
HALOE 098/0235 35916 On
The beta angle is now increasing toward a maximum angle of 45.7 degrees on
19 April 1998 (DOY 109).
A manually generated SSPP Stellar Offset experiments requested by SOLSTICE
was performed on GMT Day 104.
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 25.1 V end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery 2
and Battery 3 of 1.0 degrees C. Additional battery performance data for
each battery on 11 April 1998 and 17 April 1998 are provided in Appendix D.
The clock error ranged from -10.6 to +12.6 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
% ALERTS = none
% OUT-OF-LIMITS = 07
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT 03
% ACRIM 01
% HRDI 01
% MLS 01
% SUSIM 02
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 02
One (1) Anomaly Reports remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 98/062 - 2287.496945
Transponder B: 98/062 - 2287.497010
GROUND SYSTEM OPERATION
Two (2) operational support problem(s) occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 98/106 show the observatory data loss to be
20 hours, 42 minutes, 53.076 seconds (an increase of 16.554 seconds since
last report period). This is a 0.0360 percent data loss which equals a
99.9640 percent data capture for the mission.
The increase of 16.554 seconds occurred over a two day period as follows:
7 8.000 seconds on Day 100 due to insufficient overlap, and
7 8.544 seconds on Day 106 due to insufficient overlap.
SOLSTICE operations has submitted a requested performance of two stellar
offset experiment on GMT Day 111 and 113.
The FOT continues to experience along track ephemeris transition
comparisons in excess of 1.5 km when processing the ephemeris in the
Command Management System (CMS) software. FDF is still investigating the
possible need to transition to more frequent spacecraft ephemeris
deliveries (i.e. back to a daily delivery schedule instead of the current
Monday/Wednesday/Friday schedule).
No IP testing was performed by the UARS FOT this week. The FOT continues to
work with MSOCC personnel to finalize the details of configuring a forward
link command loop-back test using the UARS IP conversion devices (CDs)
located in the MSOCC area
The FOT has agreed to a new report format suggested by FDF that lists the
weekly UARS Transponder Frequency measurements. Previously, this data was
supplied as a CMS raw database file. The file had to be processed on the
CMS each week so that it could be converted into a readable report. Then,
the frequency information would be manually entered into a MS Excel
spreadsheet for trending purposes. Now, FDF simply e-mails a MS Word
document (which contains the same information as the old CMS file) to the
lead mission planner. No CMS processing is required to convert the
frequency data to a readable format.
The FOT continued making modifications to the OBCSTF resident POCC
database. These changes will be incorporated into the next operational
database release, version 66.
OTHER SIGNIFICANT EVENTS
Work on the upgrade of the OBC Power Monitor (PMON) software continued.
The software is completed and efforts are directed to support areas such as
database and supporting software. New RTS assignments to support SOLSTICE
day-only operation (when MLS & HALOE are on simultaneously) and enhanced
MLS and HRDI heater management have been defined. These new RTSs will be
included with the new PMON software uplink.
Work continues on the UARS August orbit adjust planning. Data on the latch
valve and piping was located and a quick evaluation indicates that a safe
refill procedure may be worked out with the help of the GSFC Propulsion
Engineering Group. The FOT is accumulating a UARS data package which will
allow an independent evaluation of and plan for the B/D thruster line
refill by the GSFC Propulsion Engineering group. However their personnel
are fully committed to other projects until May 1998, and until a Key
person (who has worked on this problem before) returns from leave.
FOT continues to support CNMOS in their efforts to make necessary software
changes (MSOCC, CMS, DCF, etc.) to support the year 2000 rollover.
Monitoring of the magnetic field model performance continues.
FOT continues to improve a Power Profile for UARS based on the current
status of the Modular Power System.
FOT continues to support the new UARS ground system re-engineering effort.
MEETINGS
No UARS Battery telecon was held this week. The next telecon has not been
scheduled.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver is currently scheduled for 05 May 1998.
A second building 23 power outage is scheduled from about 12 p.m., 24 April
1998, until about 12 p.m., 27 April 1998.
The UARS orbit altitude is dropping slowly and a drag makeup orbit adjust
is targeted for 05 August 1998.
PERSONNEL CHANGES
None.
Attachments: Appendix A, Real-time Commands Sheets
Appendix B Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 17 April 1998
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
35979 102/0810 98-129 OBC Clk Adj fr 0.00 to 20.6 msec/day
36017 104/2014 98-130 OBC Clk Adj fr 20.6 to 0.00
36019 104/2334 98-131 SUSIM Power OFF
36019 104/2334 98-132 ACRIM Power OFF
36019 104/2334 98-133 HRDI Low Power
36019 105/0015 98-134 MLS Power ON
36028 105/1520 98-135 SUSIM RTS12 Reload - SUSIM Heaters On
36045 106/1734 98-136 OBC Clk Adj fr 0.00 to 20.6 msec/day
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 17 April 1998
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
WIEUTEMP Y-LO WINDII EU Temperature 35960-36021
QTSISAIF Y-HI SSPP Baseplate by SISA I/F 35968,36023-63
HRPTSPBIMON Y-HI PTS Pulse Bus Current Monitor 35983
QTSSPPMX Y-HI SSPP -x Thermostat 36023-63
Numerous SOLSTICE temperatures due to on/off cycling
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
None this report period
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 17 April 1998
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
35979 102/08:15 (12 Apr 98) from 00.0 to 20.6 msec/day
36017 104/20:21 (14 Apr 98) from 20.6 to 00.0 msec/day
36045 106/17:38 (16 Apr 98) from 00.0 to 20.6 msec/day
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 17 April 1998
11 Apr 1998 - GMT Day 101 Beta = 32.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -2.73/-3.68 +0.0/ +0.0 18.4 0.000 0.0
2 +39.2/ -11.2 +6.44/+5.15 +30.4/-15.6 25.1 1.022 22.7
3 +67.2/ -11.2 +5.47/+4.19 +29.6/-16.4 25.1 1.026 22.2
17 Apr 1998 - GMT Day 107 Beta = 45.0 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.05/-4.31 +0.0/ +0.0 18.4 0.000 0.0
2 +44.8/ -11.2 +6.44/+5.15 +28.4/-16.0 25.1 1.022 22.4
3 +67.2/ -11.2 +5.15/+3.87 +27.2/-16.0 25.1 1.026 22.0
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes
APPENDIX E
UARS ANOMALY REPORTS
Week ending 17 April 1998
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-015 FOT 35961 101/0329 14170 Generic Late Acquisition
98-015 FOT 35983 102/1459 14170 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-017 FOT 35955 100/18:10 TAC 4 CRASH
98-018 FOT 35958 100/23:23 SOLSTICE LD PROB
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-015 Generic Late Acq for April 1998 (Open until 30 April 1998).